OpenStudy (anonymous):

Part 1: Create your own radical expression. (1 point) Part 2: Convert it to a rational exponent. (2 points) Part 3: Explain, in complete sentences, how the expression was converted. (3 points) HELP!!!!

OpenStudy (anonymous):

1. 2*sqrt(5) 2. 2*5^1/2 3. n-root(x) = x^1/n, in this case: 2*square(2)-root(5) = x^1/2

OpenStudy (anonymous):

That wasn't really complete sentences though d:

OpenStudy (anonymous):

what's that website that gives you the answers and shows you how it works stuff out??

OpenStudy (anonymous):

Do you understand how to convert??? And what website do you mean?

OpenStudy (anonymous):

idk i forgot what it started with... But everyone on here says to go to this one website

OpenStudy (anonymous):

http://www.wolframalpha.com/ perhaps?

OpenStudy (anonymous):

yes!!!

OpenStudy (anonymous):

The number 3 doesn't make sense

OpenStudy (anonymous):

3 where?

OpenStudy (anonymous):

but 1 is, 2√5 and 2 is 2 • 5 ^1/2

OpenStudy (anonymous):

Oh, alright what I meant was the n-root of "something", or x, is "something" to the power of 1/n. n-root(x) =x^1/n

OpenStudy (anonymous):

Say, the fifth-root of 415, which is the same as 415^1/5